OpenStudy (sissyedgar):

Which ordered pairs are solutions to the inequality y−2x≤−3 Select each correct answer. (5, −3) (1, −1) ♥ (−6, −3) (0, −2) ♥ (7, 12) ♥ ♥Check my answers please. Thanks. ♥

satellite73 (satellite73):

(0, −2) ♥ \[-2-2\times 0\leq -3\]?

OpenStudy (sissyedgar):

Ooh, It equals -2. Right...

satellite73 (satellite73):

(5, −3) \[-3- 2\times 5\]

OpenStudy (sissyedgar):

-8.. Well I missed a lot.

satellite73 (satellite73):

(7, 12) ♥\[12-2\times 7\]

OpenStudy (sissyedgar):

-2. Which is greater than

satellite73 (satellite73):

yeah so you missed a couple

OpenStudy (sissyedgar):

Well thanks! x)

satellite73 (satellite73):

yw